# CURE: Concept Unlearning via Orthogonal Representation Editing in Diffusion Models

### v1

@@ 8.1 Extended Inappropriate Content Removal
8.2 Additional Qualitative Results
@@
Fig. 10 provides visual samples for further assessing the removal of an artistic style, and evaluating the interference of different unlearning approaches on untargeted artistic styles. The images enclosed in red dotted borders are the intended erasure, and the off-diagonal images show effect on different styles. While SLD [16] fails to preserve the original composition in SDv1.4 (first column), marking obvious impact on general image generation capabilities of the unlearnt model, it also fails to show acceptable unlearning differnce between the images marked for intended erasure and the untargeted images. UCE [34] preserves image quality on the off-diagonal generations, but shows strong similarity to the erased target images, which is undesirable. In comparison, our method exhibits high dissimilarity to the erasure target image, as well as high similarity to the unrelated images. Hence, our approach excels in both aspects of effective unlearning and minimal impact on unintended concepts.
Fig. 11 shows images conditioned on COCO-$30$k’s captions. A good unlearning method should produce well-aligned images for unerased concepts. Methods like SLD [16] struggles to generate the a young woman carrying an umbrella, as seen in the third row.
8.3 License Information

### v2

@@ B.1 Extended Inappropriate Content Removal
B.2 Additional Qualitative Results
@@
Fig. 10 provides visual samples for further assessing the removal of an artistic style, and evaluating the interference of different unlearning approaches on untargeted artistic styles. The images enclosed in red dotted borders are the intended erasure, and the off-diagonal images show effect on different styles. While SLD [16] fails to preserve the original composition in SDv1.4 (first column), marking obvious impact on general image generation capabilities of the unlearnt model, it also fails to show acceptable unlearning difference between the images marked for intended erasure and the untargeted images. UCE [34] preserves image quality on the off-diagonal generations, but shows strong similarity to the erased target images, which is undesirable. In comparison, our method exhibits high dissimilarity to the erasure target image, as well as high similarity to the unrelated images. Hence, our approach excels in both aspects of effective unlearning and minimal impact on unintended concepts.
Fig. 11 shows images conditioned on COCO-3030k’s captions. A good unlearning method should produce well-aligned images for unerased concepts. Methods like SLD [16] struggles to generate the a young woman carrying an umbrella, as seen in the third row.
@@
Appendix C Additional Discussions
Optional Retain Set. Let $\mathcal{P}_{f}$ and $\mathcal{P}_{r}$ denote the orthogonal projectors for the forget and retain subspaces, respectively, and let $\alpha>0$ be the spectral suppression strength. CURE supports an optional retain set: when no retain supervision is provided, we set $\mathcal{P}_{r}=\mathbf{0}$ and the operator reduces to a forget-only form that still performs targeted erasure via spectral shrinkage along discriminative directions of the target concept.
While $\mathcal{P}_{r}$ offers a fine-grained control knob when available, CURE remains effective without it: in our main evaluations (Tables 1234) we set $\mathcal{P}_{r}=\mathbf{0}$ and still observe strong erasure with minimal degradation to unrelated content (Figs. 13457). This design choice reflects realistic black-box deployments where “safe” concepts are under-specified and suppression is driven instead by the spectral parameter $\alpha$, which concentrates attenuation along the most discriminative directions of the target subspace, implicitly preserving neighboring semantics. When desired, users may supply $\mathcal{P}_{r}$ for added specificity (e.g., artist erasure in Fig. 6).
Scope of Guarantees. CURE is grounded in spectral geometry and admits a Tikhonov-style interpretation (Sec A), which clarifies how forgetting pressure is distributed across subspaces. However, we do not claim formal guarantees on global optimality or worst-case trade-offs between the forget and retain subspaces. Establishing such guarantees – e.g., bounds on leakage of erased concepts under prompt perturbations while preserving retain fidelity – remains an important direction for future work.
Data/Model Dependence. Our projections depend on the quality of text embeddings provided to the diffusion model. While our experiments show strong robustness across objects, styles, identities, and NSFW concepts (Main Manuscript Tables 1234; Appendix Table 7), the framework inherits any systematic biases or brittleness from the underlying text encoder.
Setting the NudeNet Threshold. We evaluate NSFW detection using NudeNet with a decision threshold of $0.6$. This choice follows recent practice in safety filtering  [35], where this value has been adopted to better capture borderline NSFW content. This threshold ensures compatibility with safety-sensitive applications by being sufficiently conservative. Importantly, for fairness and consistency, all methods in our evaluation, including baselines, have been assessed using this same threshold.
Subspace construction and Prompt Templates. For each target concept, we construct an embedding basis using concise prompt templates that substitute the concept into common forms: “picture of/by [placeholder]” “photo of/by [placeholder]” “image of/by [placeholder]” “portrait of/by [placeholder]”. This is consistent with prior works [34, 35]. Empirically, we observe using 33-55 diverse prompts suffices to construct a stable and expressive embedding basis. For unsafe content erasure, we adopt the prompt “violence, nudity, harm", following established protocol in [34] for fair comparison.
Computational Overhead for CURE. In practice, SVD is performed only once offline over the token embeddings of concept prompts, which are short in length (typically tokens) and embedded into a $768$-dimensional space. Thus, the actual SVD computation is extremely lightweight – on the order of milliseconds on a CPU – and not a bottleneck. Once the forget/retain subspaces are computed, the weight projection is a one-time linear algebra operation, requiring under $2$ seconds on GPU for all layers. We further provide supporting details for this:
  • Token Embedding Extraction: For each concept to be forgotten (e.g., "cat"), we extract token embeddings using Stable Diffusion’s text encoder. This typically results in a small matrix of size $n_{tokens}\times 768$. For example, the prompt "cat” yields $2$ tokens, which can be expanded using related phrases like "a picture of [placeholder concept]" to around $6$ tokens. This step is equivalent to a single text encoder forward pass and takes less than $0.1$ seconds.
  • SVD Computation: We perform reduced SVD on the token matrix (e.g., $6\times 768$), retaining top-$k$ components (typically $k\leq 5$). This computation is lightweight and completes in under $0.5$ seconds on a single GPU using standard libraries like torch.linalg.svd.
  • Projection Operator Construction: We construct the projection matrix $\mathcal{P}_{dis}$ using spectral expansion with strength parameter $\alpha$ (e.g., $\alpha=2$). This matrix operation is inexpensive and completes in less than $0.1$ seconds.
  • Cross-Attention Weight Update: We apply the resulting operator to the key and value weights in the cross-attention layers of Stable Diffusion’s U-Net. Since the edit is closed-form and affects a fixed number of layers, the update takes roughly $1.2$ seconds end-to-end.
As shown in Table 5, the entire CURE operation, including all steps above, completes in under $2$ seconds on a single GPU. In contrast, methods like [27, 29] report $\sim 4500$s and $500$s respectively. CURE is thus orders of magnitude faster, requires no gradient computation, and avoids costly training pipelines. We further note that the runtime is independent of image resolution or prompt complexity, making CURE practical for scalable deployment.
Appendix D License Information